OpenStudy (anonymous):

can we use two different suppositions for a variable in two different terms of the same equation? e.g. 2*c1+3*c2+{sumation 0-inf} k*(k+1)+ {sumation 2-inf}ck*(k-1) = 0

OpenStudy (anonymous):

can we say: let k=j+2 for 3rd term and k=j for 4th term [the equation is not real]

OpenStudy (phi):

not sure what you are asking. Can you ask again in a different way, or post a specific problem you want to solve?

OpenStudy (anonymous):

in a single eq we have number of terms having a variable k, can we make supposition for k for the specific term in the eq?

OpenStudy (anonymous):

\[-c_{0}x + \sum_{k=0}^{\infty} c _{k} k(k+1)x^0 + \sum_{k=2}^{\infty} c_{k} k(k-1)x^1 = 0\] is it legal to make a supposition like, let k=j-2 for 3rd term and k=j for 4th term which will turn the equation into \[-c_{0}x + \sum_{j=2}^{\infty}c_{j-2}(j-2)(j-1)x^0 + \sum_{j=2}^{\infty}c_{j}j(j-1)x^1 = 0\] ? ?

OpenStudy (phi):

is it legal to make a supposition like, let k=j-2 for 3rd term and k=j for 4th term I would call that "renaming" the variable. you can let k= j-2 (which also means j= k+2) you will still get the same terms when you expand the summation, so it is ok to do this.

OpenStudy (anonymous):

Q: renaming the variable k=j+2 for a term and k=j for the other one in the same eq? still legal?

OpenStudy (phi):

yes. as long as you don't change the actual terms of the summation you can rename things. in other words, summing k=0 to 2 of k is the same as summing j= -2 to 0 of (j+2) in both cases, the summation is 0+1+2